OpenStudy (anonymous):

Emily and Holly are collecting dolls. They have 39 dolls between them. Write an equation showing this. Emily has 7 fewer than holly write an equation showing this. How many dolls does each girl have?

jimthompson5910 (jim_thompson5910):

Let E = # of dolls Emily has, H = # of dolls Holly has

jimthompson5910 (jim_thompson5910):

They have 39 dolls between them so if you add up their individual totals (E and H), you will get 39 this means E+H = 39

jimthompson5910 (jim_thompson5910):

Emily has 7 fewer than holly which means what?

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

still cant get it

jimthompson5910 (jim_thompson5910):

Emily has 7 fewer than holly Emily = Holly - 7 E = H - 7

jimthompson5910 (jim_thompson5910):

So your two equations are E+H = 39 E = H - 7

jimthompson5910 (jim_thompson5910):

Then you plug in E = H-7 into E+H = 39 to get E+H = 39 H - 7 + H = 39 2H - 7 = 39 2H = 39 + 7 2H = 46 H = ???

OpenStudy (anonymous):

ohh ok thanx

jimthompson5910 (jim_thompson5910):

So what do you get for E and H